# Non-uniqueness of phase transitions for graphical representations of the Ising model on tree-like graphs

### v1

@@ Abstract.
We consider the graphical representations of the Ising model on tree-like graphs. We construct a class of graphs on which the loop $\mathrm{O}(1)$ model and then single random current exhibit a non-unique phase transition with respect to the inverse temperature, highlighting the non-monotonicity of both models. It follows from the construction that there exist infinite graphs $\mathbb{G}\subseteq\mathbb{G}^{\prime}$ such that the uniform even subgraph of $\mathbb{G}^{\prime}$ percolates and the uniform even subgraph of $\mathbb{G}$ does not. We also show that on the wired $d$-regular tree, the phase transitions of the loop $\mathrm{O}(1)$, the single random current, and the random-cluster models are all unique and coincide.
@@ 1. Introduction
The Ising model needs no introduction as one of the cornerstones of statistical mechanics, and over the past 50 years its so-called graphical representations have become one of the main tools for its rigorous study [1, 4, 5]. Consequently, they are increasingly regarded as objects of independent study [3, 6, 7, 14, 16, 17]. The most prominent of these is the random-cluster model [^1] We only consider the case of cluster weight $q=2$ in which case it is also referred to as the FK-Ising model. $\phi_{x}$, introduced in [10] as an interpolation between Potts models. The loop $\mathrm{O}(1)$ model $\ell_{x}$, was introduced by Van der Waerden [25] as the high-temperature expansion of the Ising model. On finite graphs, $\ell_{x}$ can be defined as Bernoulli percolation $\mathbb{P}_{p}$ at parameter $p=\frac{x}{1+x}$ conditioned on being even (that is every vertex has even degree). The random current representation $\mathbf{P}_{x}$ was first introduced in [12] and given a useful probabilistic interpretation in [1]. While usually considered as a multigraph, we will be concerned with its traced version (i.e. the induced simple graph), which can be defined as $\mathbf{P}_{x}=\ell_{x}\cup\mathbb{P}_{1-\sqrt{1-x^{2}}}$, where $\mu_{1}\cup\mu_{2}$ is the distribution of $\omega_{1}\cup\omega_{2}$ under $\mu_{1}\otimes\mu_{2}$ ($\omega_{i}\sim\mu_{i}$). We also refer to $\mathbf{P}_{x}\cup\mathbf{P}_{x}$ as the double random current.
Similarly, we note that the random-cluster can be defined as $\phi_{x}=\ell_{x}\cup\mathbb{P}_{x}$. As all models are obtained from the loop $\mathrm{O}(1)$ model via ”sprinkling” with Bernoulli percolation, much of our analysis will focus on the former.
@@ 1.1. Results
In this paper, we prove non-uniqueness of the percolative phase transition of the (free) loop $\mathrm{O}(1)$ model $\ell_{x,\mathbb{G}}^{0}$. Here, the index $\mathbb{G}$ denotes the underlying graph and the superscript 00 denotes free boundary conditions. We also let $\mathcal{C}_{\infty}$ denote the event that there exists an infinite cluster.
@@
In Corollary 2.8 we prove the same result for the (traced, sourceless) single random current model $\mathbf{P}_{x}^{0}$. Next, consider the uniform even subgraph $\operatorname*{UEG}_{\mathbb{G}}$, defined to be the uniform measure on even subgraphs of $\mathbb{G}$. This model is intimately related to the Ising model [2, 13, 16] and can be understood as a special case of $\ell_{x,G}^{0}$ for $x=1$. Using Theorem 1.1, we prove that percolation of the uniform even subgraph is not monotone in the graph.
@@
Figure 1. The phase diagrams of the loop $\mathrm{O}(1)$, single random current, and random-cluster measures on the $d$-regular wired tree $\mathbb{T}^{d}$ coincide. The free measures on $\mathtt{C}_{n}^{d}$, the $d$-regular tree where every edge is substituted by a cycle, have different phase transitions. Finally, the free loop $\mathrm{O}(1)$ model on the monster $\mathbb{M}$ (constructed in the proof of Theorem 1.1) has a non-unique phase transition. This is to be contrasted with the corresponding table for the hypercubic and hexagonal lattices in [16, Figure 1].
@@ 1.2. The graphical representations of Ising
$\ell_{x,G}[\eta]=\frac{1}{Z_{G}}x^{\absolutevalue{\eta}},\text{ for each }\eta\in\Omega_{\emptyset}(G)$ (1)
with $Z_{G}=\sum_{\eta\in\Omega_{\emptyset}(G)}x^{\absolutevalue{\eta}}$. Here |$\absolutevalue{\eta}$ | denotes the number of (open) edges in $\eta$ and $x=\tanh(\beta)\in(0,1)$ as in [22], where $\beta$ is the inverse temperature. For $G$ a graph with boundary, we denote $\ell^{1}_{x,G}=\ell_{x,G/\sim},$ where $\sim$ identifies the boundary vertices of $G$. We refer to $\ell^{1}_{x,G}$ as the wired loop O($1$) model.
@@
where $\ell^{0}_{x}=\ell_{x}$. This definition of the model is equivalent to the standard definition due to a result by Lupu and Werner [23].
@@
which is equivalent to the standard definition of the model by a result due to Grimmet and Janson [13].
@@
We refer to the lecture notes of Duminil-Copin [5] for an overall introduction to the Ising model and its graphical representations. Furthermore, we stick with the parametrisation in terms of the loop O($1$) parameter $x\in[0,1]$ throughout this paper and refer to [16, Table 1] for an overview of the standard parametrisations.
@@
In the following, we will consider the uniform even subgraph $\operatorname*{UEG}$, which not only serves as the limiting case of the loop $O(1)$ model $\ell_{x}$ for $x=1$, but also yields (perhaps surprising) connections between the different graphical representations. For a finite graph $G,$ the uniform even subgraph $\operatorname*{UEG}_{G}$ is a uniform element of $\Omega_{\emptyset}(G)$, the set of even subgraphs of $G$. In [16], an abstract view of the uniform even subgraph was taken as the Haar measure on the group of even graphs [^3] With the group operation given by pointwise addition mod 2 in the space $\{0,1\}^{E}$ — or, equivalently, taking symmetric differences of edge sets., and its percolative properties were studied. Before that, the uniform even subgraph and its infinite volume measures were studied in detail by Angel, Ray and Spinka in [2], where the free $\operatorname*{UEG}^{0}$ and wired uniform even subgraphs $\operatorname*{UEG}^{1}$ were introduced and it was shown that they coincide on one-ended graphs [2, Lemma 3.9]. In this article, we are concerned with tree-like graphs (as opposed to, say, $\mathbb{Z}^{d}$), which in general have infinitely many ends. Hence, the distinction between the free and wired measures, $\operatorname*{UEG}^{0}$ and $\operatorname*{UEG}^{1},$ plays a bigger role than in [2, 16].
@@
The set of all finite even graphs $\Omega_{\emptyset}^{<\infty}(\mathbb{G})=\{\eta\in\Omega_{\emptyset}(\mathbb{G})\mid\absolutevalue{\eta}<\infty\}$ is a subgroup of $\Omega_{\emptyset}$. Its closure
@@
is a (compact) subgroup of $\Omega_{\emptyset}(\mathbb{G})$, and the free uniform even subgraph $\operatorname*{UEG}_{\mathbb{G}}^{0}$ is the Haar measure on that group. For more details on the construction of the free and wired uniform even subgraphs for infinite graphs see [16, Section 3.2].
In [13, Theorem 3.5] it was realised that the loop $\mathrm{O}(1)$ model arises as the uniform even subgraph of the random-cluster model and in [19, Theorem 4.1], that it is also the uniform even subgraph of the double random current.
@@
where $\xi=0$ in the free case and $\xi=1$ in the wired case [^4] For finite graphs we write $\ell$, omitting the boundary condition., and $\phi^{\xi}_{\mathbb{G}}$ is defined as a thermodynamic limit (see e.g. [16, Sec. 2.1.3.]) when $\mathbb{G}$ is infinite. In infinite volume, this may be taken as the definition of the loop O($1$) model (cf. [16, (4)]). This can then be used to define the single and double random current,
@@ 2. Non-uniqueness of percolation
Let $G\subseteq G^{\prime}$ be two infinite graphs and suppose that the uniform even subgraph of $G^{\prime}$ almost surely percolates. Does the uniform even subgraph of $G$ almost surely percolate?
@@
We use the fact that the loop $O(1)$ model factorises on graphs lacking certain cycles: A cycle denotes a path of vertices $v_{1},v_{2},...,v_{n}$ such that $v_{1}=v_{n}$. We say that a cycle is simple if $v_{j}\neq v_{k}$ for distinct $1\leq j,k<n$.
@@
There exist parameter values $10<x_{1}<x_{2}<$ 1, and a finite graph $G^{\diamond}$ with vertices $a,b$ such that
@@
Figure 2. The graph $G^{\diamond}$ (pictured to the right) along with its eight even subgraphs (including $G^{\diamond}$ itself). We let the outer paths be $n$ edges long and the inner paths be $m$ edges long. The nodes $a$ and $b$ are marked with dots. We list the number of edges of each subgraph, the corresponding weights and whether $a$ and $b$ are connected in the subgraph. (Sketch and text partially revised from [19]
Subgraph .. .. .. .. . . .. .. ..$m$$m$$n$$n$
@@
Figure 3. Graph of the connection probability for the loop $\mathrm{O}(1)$ model on the graph described in Figure 2 for $n=12$ and $m=2$. See also [19, Figure 2.3] for similar figures.
@@
We construct $\mathbb{M}$ from the $d$-regular tree $\mathbb{T}^{d}$ with root 00 for an appropriate choice of $d$. Consider the natural orientation of the tree where edges are oriented away from the root and replace every such oriented edge $e=(v,w)$ by a copy of the graph $G^{\diamond}$ where $a$ is identified with $v$ and $b$ is identified with $w$. More formally, let $\mathbb{M}=(\coprod_{e\in E(\mathbb{T}^{d})}G_{e}^{\diamond})/{\sim},$ where $\sim$ is the equivalence relation such that $a_{e}\sim b_{e^{\prime}}$ whenever the source of $e$ is equal to the sink of $e^{\prime}$. See Figure 4 for an illustration. Now, since the macroscopic structure of $\mathbb{M}$ is that of a tree, Lemma 2.4 applies, and it holds that
@@
[…]
@@
Figure 4. An example of the graph $\mathbb{M}$ built from the graphs $G^{\diamond}$ (see Figure 2) when $d=2$. To the left with the $d$-regular tree overlaid. To the right, the geometry of $\mathbb{M}$ at a single vertex of the initial tree.
@@
In this section, we will construct a supergraph [^5] One may note that our trick is basically to pick a one-ended supergraph of a two-ended graph. One may check that the wired UEG of a graph with multiple ends percolates (see also [16, Eq. (14)]). Therefore, our construction does not have the flavour of an optimal solution to the problem. of $\mathbb{Z}$ for which the wired UEG does not percolate.
Define $\mathbb{G}=(\mathbb{Z},\hat{\mathbb{E}}),$ where $(n,m)\in\hat{E}$ if either $|n-m|=1$ or $n=-m$. Let $\hat{\mathbb{E}}_{\text{arc}}:=\hat{\mathbb{E}}\setminus\mathbb{E}(\mathbb{Z})$ and $\mathbb{G}_{\text{arc}}=(\mathbb{Z},\hat{E}_{\text{arc}})$.
@@
This follows from the fact that the restriction group homomorphism $\Omega_{\emptyset}(G)\to\{0,1\}^{\hat{\mathbb{E}}_{\text{arc}}}$ is surjective. To see that the map is surjective, it suffices that all single-edge configurations lie in the image. For a given $e=(-n,n)\in\hat{\mathbb{E}}_{\text{arc}},$ we see that $1\!\!1_{e}$ is the image of the loop containing $[-n,n]\cap\mathbb{Z}$ and $e$. ∎
Proof of Corollary 1.2 (for the wired case). Let $\mathbb{G}$ be as above. Let $v\in\mathbb{Z}$ and note that on the event that there is an open edge in $\hat{\mathbb{E}}_{\text{arc}}$ outside of $[-|v|,|v|],$ the cluster of $v$ has to be finite. By Lemma 2.6, with probability 1, infinitely many edges in $\hat{\mathbb{E}}_{\text{arc}}$ are open in $\operatorname*{UEG}_{\mathbb{G}}$. By the previous comment, on this event, all clusters are finite. Hence $\operatorname*{UEG}_{\mathbb{G}}$ does not percolate. However, $\mathbb{Z}$ has exactly two even subgraphs, $\eta\equiv 0$ and $\eta\equiv 1$, so $\operatorname*{UEG}_{\mathbb{Z}}^{1}$ does percolate. ∎

### v2

@@ Abstract.
We consider the graphical representations of the Ising model on tree-like graphs. We construct a class of graphs on which the loop $\mathrm{O}(1)$ model and the single random current exhibit a non-unique phase transition with respect to the inverse temperature, highlighting the non-monotonicity of both models. It follows from the construction that there exist infinite graphs $\mathbb{G}\subseteq\mathbb{G}^{\prime}$ such that the uniform even subgraph of $\mathbb{G}^{\prime}$ percolates and the uniform even subgraph of $\mathbb{G}$ does not. We also show that on the wired $d$-regular tree, the phase transitions of the loop $\mathrm{O}(1)$, the single random current, and the random-cluster models are all unique and coincide.
@@ 1. Introduction
The Ising model needs no introduction as one of the cornerstones of statistical mechanics, and over the past 50 years its so-called graphical representations have become one of the main tools for its rigorous study [1, 4, 5]. Consequently, they are increasingly regarded as objects of independent interest [3, 8, 7, 16, 18, 19]. The most prominent of these is the random-cluster model [^1] We only consider the case of cluster weight $q=2$ in which case it is also referred to as the FK-Ising model. $\phi_{x}$, introduced in [12]. The loop $\mathrm{O}(1)$ model $\ell_{x}$, was introduced by Van der Waerden [28] as the high-temperature expansion of the Ising model. On finite graphs, $\ell_{x}$ can be defined as Bernoulli percolation $\mathbb{P}_{p}$ at parameter $p=\frac{x}{1+x}$ conditioned on being even (that is every vertex has even degree). The random current representation $\mathbf{P}_{x}$ was first introduced in [14] and given a useful probabilistic interpretation in [1]. While usually considered as a multigraph, we will be concerned with its traced version (i.e. the induced simple graph), which can be defined as $\mathbf{P}_{x}=\ell_{x}\cup\mathbb{P}_{1-\sqrt{1-x^{2}}}$, where $\mu_{1}\cup\mu_{2}$ is the distribution of $\omega_{1}\cup\omega_{2}$ under $\mu_{1}\otimes\mu_{2}$ ($\omega_{i}\sim\mu_{i}$). We also refer to $\mathbf{P}_{x}\cup\mathbf{P}_{x}$ as the double random current.
Similarly, we note that the random-cluster can be defined as $\phi_{x}=\ell_{x}\cup\mathbb{P}_{x}$ by [11, 15]. As all models are obtained from the loop $\mathrm{O}(1)$ model via ”sprinkling” with Bernoulli percolation, much of our analysis will focus on the former.
@@ 1.1. Results
In this paper, we prove non-uniqueness of the percolative phase transition of the (free) loop $\mathrm{O}(1)$ model $\ell_{x,\mathbb{G}}^{0}$ (the infinite volume measure is defined in (4) below). Here, the index $\mathbb{G}$ denotes the underlying graph and the superscript 0 denotes free boundary conditions. We also let $\mathcal{C}_{\infty}$ denote the event that there exists an infinite cluster.
@@
In Corollary 2.8 we prove the same result for the (traced, sourceless) single random current model $\mathbf{P}_{x}^{0}$. Next, consider the uniform even subgraph $\operatorname*{UEG}_{\mathbb{G}}$, defined to be the uniform measure on even subgraphs of $\mathbb{G}$. This model is intimately related to the Ising model [2, 15, 18] and can be understood as a special case of $\ell_{x,G}^{0}$ for $x=1$. Using Theorem 1.1, we prove that percolation of the uniform even subgraph is not monotone in the graph. The result holds for both the free and wired infinite volume measures, both of which will be defined below.
@@
Figure 1. The phase diagrams of the loop $\mathrm{O}(1)$, single random current, and random-cluster measures on the $d$-regular wired tree $\mathbb{T}^{d}$ coincide. The free measures on $\mathtt{C}_{n}^{d}$, the $d$-regular tree where every edge is substituted by a cycle, have different phase transitions. Finally, the free loop $\mathrm{O}(1)$ model on the monster $\mathbb{M}$ (constructed in the proof of Theorem 1.1) has a non-unique phase transition. This is to be contrasted with the corresponding table for the hypercubic and hexagonal lattices in [18, Figure 1].
@@ 1.2. The graphical representations of Ising
[…]
with $|Z_{G}=\sum_{\eta\in\Omega_{\emptyset}(G)}x^{\absolutevalue{\eta}}$. Here $|\absolutevalue{\eta}$ denotes the number of (open) edges in $\eta$ and $x=\tanh(\beta)\in(0,1)$ as in [25], where $\beta$ is the inverse temperature. For $G$ a graph with boundary, we denote $\ell^{1}_{x,G}=\ell_{x,G/\sim},$ where $\sim$ identifies the boundary vertices of $G$. We refer to $\ell^{1}_{x,G}$ as the wired loop O($1$) model.
@@
where $\ell^{0}_{x}=\ell_{x}$. This definition of the model is equivalent to the standard definition as explained in [6, Remark 3.4], see also [26].
@@
which is equivalent to the standard definition of the model by the results [11, 15].
@@
We refer to the lecture notes of Duminil-Copin [5] for an overall introduction to the Ising model and its graphical representations. Furthermore, we stick with the parametrisation in terms of the loop O($1$) parameter $x\in[0,1]$ throughout this paper and refer to [18, Table 1] for an overview of the standard parametrisations.
@@
In the following, we will consider the uniform even subgraph $\operatorname*{UEG}$, which not only serves as the limiting case of the loop $O(1)$ model $\ell_{x}$ for $x=1$, but also yields (perhaps surprising) connections between the different graphical representations. For a finite graph $G,$ the uniform even subgraph $\operatorname*{UEG}_{G}$ is a uniform element of $\Omega_{\emptyset}(G)$, the set of even subgraphs of $G$. In [18], an abstract view of the uniform even subgraph was taken as the Haar measure on the group of even graphs [^3] With the group operation given by pointwise addition mod 2 in the space $\{0,1\}^{E}$ - or, equivalently, taking symmetric differences of edge sets., and its percolative properties were studied. Before that, the uniform even subgraph and its infinite volume measures were studied in detail by Angel, Ray and Spinka in [2], where the free $\operatorname*{UEG}^{0}$ and wired uniform even subgraphs $\operatorname*{UEG}^{1}$ were introduced and it was shown that they coincide on one-ended graphs [2, Lemma 3.9]. In this article, we are concerned with tree-like graphs (as opposed to, say, $\mathbb{Z}^{d}$), which in general have infinitely many ends. Hence, the distinction between the free and wired measures, $\operatorname*{UEG}^{0}$ and $\operatorname*{UEG}^{1},$ plays a bigger role than in [2, 18].
@@
[…]
@@
is a (compact) subgroup of $\Omega_{\emptyset}(\mathbb{G})$, and the free uniform even subgraph $\operatorname*{UEG}_{\mathbb{G}}^{0}$ is defined to be the Haar measure on that group. For more details on the construction of the free and wired uniform even subgraphs for infinite graphs and a characterization of the Gibbs of the uniform even graph on trees see [18, Section 3.2].
In [15, Theorem 3.5] it was realised that the loop $\mathrm{O}(1)$ model arises as the uniform even subgraph of the random-cluster model. Similarly, loop $\mathrm{O}(1)$ model arises as the uniform even subgraph of the double random current which appeared implicitly in [24, pf. of Theorem 3.2] and was proven in [21, Theorem 4.1].
@@
where $\xi=0$ in the free case and $\xi=1$ in the wired case [^4] For finite graphs we write $\ell$, omitting the boundary condition., and $\phi^{\xi}_{\mathbb{G}}$ is defined as a thermodynamic limit (see e.g. [18, Sec. 2.1.3.]) when $\mathbb{G}$ is infinite. In infinite volume, this may be taken as the definition of the loop O($1$) model (cf. [18, (4)]). This can then be used to define the single and double random current,
@@ 2. Non-uniqueness of percolation
Let $G\subseteq G^{\prime}$ be two infinite graphs and suppose that the uniform even subgraph of $G$ almost surely percolates. Does the uniform even subgraph of $G^{\prime}$ almost surely percolate?
@@
We use that the loop $O(1)$ model factorises on graphs lacking certain cycles: A cycle denotes a path of vertices $v_{1},v_{2},...,v_{n}$ such that $v_{1}=v_{n}$. We say that a cycle is simple if $v_{j}\neq v_{k}$ for distinct $1\leq j,k<n$.
@@
In Lemma 2.4 we proved the cut-point factorisation property for finite graphs. The infinite volume measures also have the cut-point factorisation property as they are either limits or (sprinkled) uniform even subgraphs of a cut-point factorising measure.
@@
There exist parameter values $0<x_{1}<x_{2}<1$, and a finite graph $G^{\diamond}$ with vertices $a,b$ such that
@@
Figure 2. The graph $G^{\diamond}$ (pictured to the right) along with its eight even subgraphs (including $G^{\diamond}$ itself). We let the outer paths be $n$ edges long and the inner paths be $m$ edges long. The nodes $a$ and $b$ are marked with dots. We list the number of edges of each subgraph, the corresponding weights and whether $a$ and $b$ are connected in the subgraph. (Sketch and text partially revised from [21].)
Subgraph .. .. .. .. . . .. .. ..mmmmnnnn
@@
Figure 3. Graph of the connection probability for the loop $\mathrm{O}(1)$ model on the graph $G^{\diamond}$, described in Figure 2, for $n=12$ and $m=2$. See also [21, Figure 2.3] for similar figures.
@@
We construct $\mathbb{M}$ from the $d$-regular tree $\mathbb{T}^{d}$ with root 0 for an appropriate choice of $d$. Consider the natural orientation of the tree where edges are oriented away from the root and replace every such oriented edge $e=(v,w)$ by a copy of the graph $G^{\diamond}$ where $a$ is identified with $v$ and $b$ is identified with $w$. More formally, let $\mathbb{M}=(\coprod_{e\in E(\mathbb{T}^{d})}G_{e}^{\diamond})/{\sim},$ where $\sim$ is the equivalence relation such that $a_{e}\sim b_{e^{\prime}}$ whenever the source of $e$ is equal to the sink of $e^{\prime}$. See Figure 4 for an illustration. Now, since the macroscopic structure of $\mathbb{M}$ is that of a tree, Lemma 2.4 applies, and it holds that
@@
As a consequence, analysing percolation of $\ell_{x,\mathbb{M}}$ just boils down to Bernoulli percolation on $\mathbb{T}^{d}$: For given $\eta\in\Omega^{0}_{\emptyset}(\mathbb{M})$ and $e\in E(\mathbb{T}^{d}),$ we define $m_{e}=1$ if $a\overset{\eta}{\leftrightarrow}b$ in $G^{\diamond}_{e}$ and $m_{e}=0$ otherwise. In other words, $m=(m_{e})_{e\in E(\mathbb{T}^{d})}$ maps a percolation configuration in $\mathbb{M}$ onto one in $\mathbb{T}^{d}$ such that $0\,\smash{\overset{m(\eta)}{\longleftrightarrow}}\,\infty$ if and only if $0\,\smash{\overset{\eta}{\longleftrightarrow}}\,\infty$. Furthermore, by (9), the image measure of $\ell_{x,\mathbb{M}}$ is Bernoulli percolation: $m(\ell^{0}_{x,\mathbb{M}})=\mathbb{P}_{f(x),\mathbb{T}^{d}}$ where $f(x)=\ell_{x,G^{\diamond}}[a\leftrightarrow b]$.
@@
Figure 4. An illustration of a part of the graph $\mathbb{M}$ built from the graphs $G^{\diamond}$ (see Figure 2) when $d=2$. To the left with the $d$-regular tree overlaid. To the right, the geometry of $\mathbb{M}$ at a single vertex of the initial tree.
@@
In this section, we will construct a supergraph [^5] One may note that our trick is basically to pick a one-ended supergraph of a two-ended graph. One may check that the wired UEG of a graph with multiple ends percolates (see also [18, Eq. (14)]). Therefore, our construction does not have the flavour of an optimal solution to the problem. of $\mathbb{Z}$ for which the wired UEG does not percolate.
[…]
@@
This follows from the fact that the restriction group homomorphism $\Omega_{\emptyset}(\mathbb{G})\to\{0,1\}^{\hat{\mathbb{E}}_{\text{arc}}}$ is surjective. To see that the map is surjective, it suffices that all single-edge configurations lie in the image. For a given $e=(-n,n)\in\hat{\mathbb{E}}_{\text{arc}},$ we see that $1\!\!1_{e}$ is the image of the loop containing $n,n]\cap\mathbb{Z}$ and $e$. ∎
Proof of Corollary 1.2 (for the wired case). Let $\mathbb{G}$ be as above. Let $v\in\mathbb{Z}$ and note that on the event that there is an open edge in $\hat{\mathbb{E}}_{\text{arc}}$ outside of [|v|,|v|],[-|v|,|v|], the cluster of $v$ has to be finite. By Lemma 2.6, with probability 1, infinitely many edges in $\hat{\mathbb{E}}_{\text{arc}}$ are open in $\operatorname*{UEG}_{\mathbb{G}}$. By the previous comment, on this event, all clusters are finite. Hence $\operatorname*{UEG}_{\mathbb{G}}$ does not percolate. However, $\mathbb{Z}$ has exactly two even subgraphs, $\eta\equiv 0$ and $\eta\equiv 1$, so $\operatorname*{UEG}_{\mathbb{Z}}^{1}$ does percolate. ∎